OpenStudy (anonymous):

How can I find the limit for n, belonging to the naturals. Of: ( sqrt(n+1) - sqrt(n) ) * sqrt(n+0.5)

OpenStudy (anonymous):

\[( \sqrt(n+1) - \sqrt(n) ) * \sqrt(n + \frac{ 1 }{ 2 })\]